OpenStudy (anonymous):

Jesse’s parents bought a new car. After owning it for two years, the car’s value was $14,792. After owning it for six years, Jesse’s parents sold it to him for $5,090. Make a graph, where x is the number of years they owned the car and y is the car’s value. Graph the car’s value at 2 years and 6 years. Determine the approximate purchase price of the car A. $17,000 B. $20,000 C. $22,000 D. $24,000

OpenStudy (nurali):

Assuming this is a linear relationship (steady, constant), then just figure out the slope: (Y2 - Y1)/(X2 - X1) = (5090 - 14792)/(6-2) = -9702/4 = -2426/1 or simply -2426 this means that every year the car's values drops $2426. now just backtrack this: if it was worth $14,792 after two years, then it must have been worth $17,218 after just one year ($14,792 + $2426). Do this again and that tells you that when it came out of the retailers it was worth $22,000.